# Right Reward Right Time for Federated Learning

### v1

@@ 3.3 R3T Incentivization under Incomplete Information Case
As such, if the IR constraint of type-$1$ client is binding, that of all other types automatically holds. This completes the proof. ∎
The IR constraints in (7) can be replaced by Lemma (3), which indicates that this type-$1$ client gains or loses nothing in participating in the training process, while other client’s utilities are higher than that of the binding one. Next, we refer to [51] to introduce the following definitions for simplifying IC constraints:
@@
1) Downward incentive constraints (DICs) are IC constraints between client types $k$ and ${k^{\prime}}$, where $11\leq{k^{\prime}}\leq k-$ 1, 2) local downward incentive constraints (LDIC) is an IC constraint between client’s types $k$ and ${k^{\prime}},\text{where }{k^{\prime}}=k-1$, 3) upward incentive constraints (UICs) are IC constraints between client’s types $k$ and ${k^{\prime}},\text{where }{k}+1\leq k^{\prime}\leq{K}$, and 4) local upward incentive constraints (LUIC) is an IC constraint between client’s types $k$ and ${k^{\prime}},\text{where }k^{\prime}={k}+1$.
@@
Therefore, we conclude that LDIC has transitivity, i.e., if there is an LDIC between type-$(k-1)$ and type-$k$ clients, then we can further extend incentive constraints from type-$(k-1)$ client to type-$1$ client, that is, all DICs hold, i.e.,
@@
which can replace IC constraints in (8). This completes the proof. ∎
Input: A set of clients $\mathcal{N}$, number of training round $T$, learning rate $\eta$, local loss function $F$, initial global parameter $w^{0}$, bonus unit coefficient $\vartheta$.
@@
[…]
@@
[…]
@@ 13                  $w^{t}_{k}\leftarrow w^{t-1}_{k}-\eta\triangledown{F}_{k}(w^{t-1}_{k},\mathcal{D}_{k})$
[…]
@@
[…]
@@ 26      $w^{t}\leftarrow\sum_{k\in\mathcal{K}}{\frac{e^{*}_{k}}{\cup_{k\in\mathcal{K}}e^{*}_{k}}}w^{t}_{k}$
[…]
@@
[…]
@@
Note that at $i=1$, we have $h(t_{0})=0$ because $t_{0}$ is the time before the training process starts. This completes the proof. ∎
@@
According to (1), $h(t_{k})=1$ and the impact of efforts in non-CLPs is not of importance as CLPs, thereby $h(t_{k^{\prime}})=h(t_{k})=1,\forall{t_{k},t_{k^{\prime}}\notin\text{CLPs}}$ This leads to $e_{k}>e_{k^{\prime}}$ based on (11). Moreover, via (3.3) and $h(t_{k^{\prime}})=h(t_{k})$, we can conclude that $r_{k}=r_{k^{\prime}}=r_{1}=\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}$. Next, with the proved conditions (i.e., $h(t_{k})=h(t_{k^{\prime}})\text{ and }r_{k}=r_{k^{\prime}}$), and the expression (1), we have $R_{k}>R_{k^{\prime}}$. This completes the proof. ∎
@@
The R3T’s game and training processes in $T$ training rounds are depicted in Algorithm 1. We have a tuple of $(e^{*}_{k},t^{*}_{k},r^{*}_{k})$ being the optimal strategy of the type-$k$ client (line 1). CLPs can be identified by using federated gradient norm metric (line 1) [11]. To address the issue of insufficient client participation and its detrimental effect on final model performance, R3T increases the number of clients during the early training phase (line 1) by offering an attractive contract set. Besides, to reduce communication overhead and budget exhaustion, R3T decreases the number of clients in non-CLPs (line 1).
4 Experimental Results
@@
4.1 Experiment Setup
Simulation parameter settings. Without loss of generality, our experimental settings refer to the setups in [19, 52]. The key parameters used in the simulations are shown in Table I.
Proof of concept. We develop a system to implement a proof of concept of R3T, leveraging smart contracts on an Ethereum blockchain using Ganache [53]. The Flower framework [54] is integrated with the blockchain to facilitate federated learning experiments for model training and evaluation. The system is deployed on an Ubuntu 22.04 system equipped with an Intel Core i7-1365U CPU@1.80 GHz and 32GB of memory. To enable interaction between clients, the cloud server, and the blockchain smart contracts, we utilize the Web3 API [55] and FastAPI frameworks [56]. Also, we develop two smart contracts for federated training and storage, and reward calculation using Solidity programming language.
Model and Datasets: We use a deep neural network with four layers comprising three convolutional neural layers with dropout to prevent overfitting, batch normalization, and max pooling for reducing spatial dimensions, and one dense layer. For model training and testing, CIFAR-10 [57] and Fashion-MNIST [58] datasets are used in a non-identical and independent (non-IID) setting. We consider a highly heterogeneous setting in which the number of data samples and data labels among clients are unbalanced by sampling $p\sim\textit{Dir}{(\alpha)}$, where $\alpha=\{0.1,0.4,0.7\}$ is the parameter of Dirichlet distribution. Besides, the learning rate $\eta=$ is 3e-4 and the batch size is 32.
@@ 4.1 Experiment Setup
  • Contract theory Without considering Time under CIC (CTWT-CIC). Time property is not considered in the existing work (e.g., [51, 31, 19, 23]), that is, $t$ is excluded from the contract item. All training time slots are treated equally.
@@
  • Linear Pricing. Under IIC, the cloud server determines a unit price $C$ for the client’s effort $e$ [21].
  • Conventional Federated Learning (CFL). FL methods integrating conventional incentive mechanisms (e.g., [23, 31, 32]), which do not consider CLPs, are used to compare with R3T proof of concept.
@@ 4.2 Contract Feasibility
Impacts of different unit effort costs. In Figs.  4(4(a))-(4(f)), under CIC and IIC in CLPs [^9] We focus on one global training round in CLPs here, but analogous trends observed in non-CLPs., the client’s fixed salary, bonus, effort, total reward, utility, and cloud utility have shown decreasing trends per contract item along with the increase of unit effort cost $\delta$, where we set $\delta=\{0.2,0.6,1\}$. In other words, clients are not willing to join and contribute efforts when $\delta$ becomes large. For example, given the tenth contract item under the IIC cases, the values of the client’s effort units are around {9.60, 3.20, 1.92} corresponding to $\delta=\{0.2,0.6,1\}$. Besides, under CIC, where the cloud server possesses complete client information, the cloud can fine-tune and optimize its contract items to incentivize clients to join early and contribute more effort. This results in higher cloud utility while offering clients higher fixed salaries and bonuses, compared to IIC (e.g., 232.28 vs. 211.13, 5.96 vs. 1.01, and 119.21 vs. 18.47 given the tenth contract item with $\delta=0.2$, respectively). In addition, due to being aware of client information, the cloud server exploits the client’s efforts as much as possible while guaranteeing IR constraints (c.f. Theorem 1) in CIC, leading client utilities over contract items equal to zero (c.f., Fig. 4(4(e))).
@@
Impacts of different cloud budgets. Figs.  5(5(a))-(5(b)) show the effect of varying cloud budgets $P$ on the cloud utility contributed by type-$k$ clients (i.e., $k$-th contract items) in a global training round under both CIC and IIC in CLPs. As $P$ increases, more clients can be incentivized to participate in and allocate efforts, resulting in greater total cloud utility over all joined clients. For instance, at $P=60$, the total cloud utility of all type-$k$ clients reaches approximately 375 in CIC and 312 in IIC, compared to 266 in CIC and 311 in IIC at $P=30$, respectively. Especially at low budgets (i.e., $P=\{10,20,30\}$), the total cloud utility under IIC exceeds that under CIC. This is supported by Assumption 1 and the fact that the cloud server is aware of clients’ types in CIC and only the distribution of clients’ types in IIC, thereby different optimal contract designs. Consequently, in CIC, the cloud server may adopt a greedy allocation strategy to exhaust efforts from higher-type clients (c.f. Theorem 1), but due to a limited budget, it leads to fewer participating clients and thus lower total cloud utility compared with that of IIC. Conversely, when the budget is huge enough (i.e., $P\geq 40$), the total cloud utility in CIC is far larger than that in IIC (i.e., 375 versus 311 at $P=60$).
@@ 4.4 Proof of Concept of R3T
Convergence rate. As illustrated in Fig. 12, R3T demonstrates its efficiency by surpassing CFL in reducing convergence rate under non-IID settings for both CIFAR-10 and Fashion-MNIST datasets. Particularly, R3T achieves a significant increase during the initial FL training phase. This improvement is attributed to its strategy of attracting the highest-quality clients by offering an optimal set of contract items, which includes timely and substantial rewards, early in the training process. For example, R3T tends to reach the final target accuracy in just 20 rounds, compared with approximately 61 rounds for CFL on the datasets.
Long-term performance. For different values of the parameter of Dirichlet distribution $\alpha=\{0.4,0.7\}$ in both datasets, both R3T and CFL maintain stable test accuracies after convergence, even though R3T selects a smaller number of participating clients in these periods (c.f., in line 1 of Algorithm 1). Under a highly non-IID setting (i.e., $\alpha=0.1$), R3T demonstrates more sustainable performance compared to CFL. In CFL, the practice of treating clients’ contributions and efforts equally across all training rounds, especially under a significant quantity and label distribution skew, detrimentally affects the long-term performance of the FL model, regardless of subsequent efforts. Besides, although R3T shows a slight drop in accuracy after CLPs due to the reduced number of selected clients, the overall test accuracy remains stable around the highest bound. These observations highlight the significant impact of attracting high-quality clients early, saving around 200%-300% in the total training time and 5.2%-47.6% in the total number of clients to achieve a close targeted learning performance.
5 Conclusion
In this work, we studied a time-aware incentive mechanism implemented on blockchain smart contracts for FL, referred to as R3T. By characterizing the interactions between rational client’s time and system capabilities, efforts, joining time, reward, and their impact on cloud utility, we proposed a contract design that incorporates the temporal factor and then derived the optimal incentive strategies for the cloud. R3T transparently provides the right reward at the right time, thereby attracting the highest-quality contributions in the early training phase, which is crucial in determining the final model performance under information symmetry and information asymmetry conditions. Simulation and proof-of-concept experimental results demonstrate the superiority of R3T over incentive mechanism benchmarks in improving cloud utility and economic efficiency, reducing convergence time and the total number of participating clients, while achieving comparable model performance.
In future work, we will consider impacts of free-rider issues and malicious attacks on the efficiency of R3T, especially in CLPs. Besides, it would be interesting to evaluate the efficiency and scalability of R3T on different blockchain platforms integrating FL systems, where the entire system is operated without relying on the cloud server.

### v2

@@ III-C R3T Incentivization under Incomplete Information Case
As such, if the IR constraint of type-$1$ client is binding, that of all other types holds. This completes the proof. ∎
The IR constraints in (7) can be replaced by Lemma (3), which indicates that this type-$1$ client gains or loses nothing in participating in training, while other client’s utilities are higher than that of the binding one. Next, we refer to [zhang2015contract] to introduce the following definitions for simplifying IC constraints:
@@
1) Downward incentive constraints (DICs) are IC constraints between client types $k$ and $k^{\prime}$}, where $1\leq{k^{\prime}}\leq k-1$, 2) local downward incentive constraints (LDIC) is an IC constraint between client’s types $k$ and ${k^{\prime}},\text{where }{k^{\prime}}=k-1$, 3) upward incentive constraints (UICs) are IC constraints between client’s types $k$ and ${k^{\prime}},\text{where }{k}+1\leq k^{\prime}\leq{K}$, and 4) local upward incentive constraints (LUIC) is an IC constraint between client’s types $k$ and ${k^{\prime}},\text{where }k^{\prime}={k}+1$.
@@
Therefore, we conclude that LDIC has transitivity, i.e., if there is an LDIC between type-(k1)(k-1) and type-$k$ clients, then we can further extend incentive constraints from type-(k1)(k-1) client to type-$1$ client, that is, all DICs hold, i.e.,
@@
which replace IC constraints in (8). This completes proof. ∎
Input: A set of clients $\mathcal{N}$, number of training rounds $T$, learning rate $\eta$, local loss function $F$, initial global parameter $w^{0}$, bonus unit coefficient $\vartheta$.
@@
3   $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\rhd{\textit{ Cloud performs:}}$}
@@
7  $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\rhd{\textit{ Clients perform:}}$}
@@ 13       $w^{t}_{k}\leftarrow w^{t-1}_{k}-\eta\triangledown{F}_{k}(w^{t-1}_{k},\mathcal{D}_{k})$
15  $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\rhd{\textit{ Blockchain performs:}}$}
@@
17  $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\rhd{\textit{ Cloud performs:}}$}
@@ 26 $w^{t}\leftarrow\sum_{k\in\mathcal{K}}{\frac{e^{*}_{k}}{\cup_{k\in\mathcal{K}}e^{*}_{k}}}w^{t}_{k}$
27  $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\rhd{\textit{ Blockchain performs:}}$}
@@
29 $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\rhd{\textit{ Blockchain performs:}}$}
@@
At $i=1$, we have $h(t_{0})=0$ because $t_{0}$ is the time before the training process starts. This completes the proof. ∎
@@
 ${\quad\quad-\text{ }\frac{\theta^{2}_{k}h^{2}(t_{k})}{\delta}\Big]},$ (38a)
@@
According to (1), $h(t_{k})=1$ and the impact of efforts in non-CLPs is not of importance as CLPs, thereby $h(t_{k^{\prime}})=h(t_{k})=1,\forall{t_{k},t_{k^{\prime}}\notin\text{CLPs}}$ . This leads to $e_{k}>e_{k^{\prime}}$ based on (11). Moreover, via (III-C) and $h(t_{k^{\prime}})=h(t_{k})$, we can conclude that $r_{k}=r_{k^{\prime}}=r_{1}=\frac{\theta^{2}_{1}h^{2}(t_{1})}{2\delta(\beta-1)}$. Next, with the proved conditions (i.e., $h(t_{k})=h(t_{k^{\prime}})\text{ and }r_{k}=r_{k^{\prime}}$), and the expression (1), we have $R_{k}>R_{k^{\prime}}$. This completes the proof. ∎
@@
The R3T’s game and training processes in $T$ training rounds are depicted in Algorithm 1. We have a tuple of $(e^{*}_{k},t^{*}_{k},r^{*}_{k})$ being the optimal strategy of the type-$k$ client (line 1). CLPs can be identified by using federated gradient norm metric (line 1) [yan2023criticalfl]. To address the issue of insufficient client participation and its detrimental effect on final model performance, R3T increases the number of clients during the early training phase (line 1) by offering an attractive contract set. Besides, to reduce communication overhead and budget exhaustion, R3T decreases the number of clients in non-CLPs (line 1).
IV Experimental Results
@@
IV-A Experiment Setup
Simulation parameter settings. Without loss of generality, our experimental settings refer to the setups in [zhan2020learning, zhao2021incentive]. We also define the function $g(\cdot)$ by $g(\cdot)=h(t)e$. The key parameters used in the simulations are shown in Table I.
Proof of concept. We develop a system to implement a proof of concept of R3T, leveraging smart contracts on an Ethereum blockchain using Ganache [ganache]. The Flower framework [beutel2020flower] is integrated with the blockchain to facilitate federated learning experiments for model training and evaluation. The system is deployed on an Ubuntu 22.04 system equipped with an Intel Core i7-1365U CPU@1.80 GHz and 32GB of memory. To enable interaction between clients, the cloud server, and the smart contracts, we utilize the Web3 API [web3api] and FastAPI frameworks [fastapi]. Also, we develop two smart contracts for federated training and storage, and reward calculation using Solidity programming language.
Model and Datasets: We use a deep neural network with four layers comprising three convolutional neural layers with dropout to prevent overfitting, batch normalization, and max pooling for reducing spatial dimensions, and one dense layer. For model training and testing, CIFAR-10 [krizhevsky2009learning] and Fashion-MNIST [xiao2017fashion] datasets are used in a non-independent and identically distributed (non-IID) setting. We consider a highly heterogeneous setting in which the number of data samples and data labels among clients are unbalanced by sampling $p\sim\textit{Dir}{(\alpha)}$, where $\alpha=\{0.1,0.4,0.7\}$ is the parameter of Dirichlet distribution. Besides, the learning rate $\eta=$ is 3e-4 and the batch size is 32.
@@ IV-A Experiment Setup
  • Contract theory Without considering Time under CIC (CTWT-CIC). Time property is not considered in the existing work (e.g., [zhang2015contract, sun2021pain, zhan2020learning, wang2022infedge]), that is, $t$ is excluded from the contract item. All training time slots are treated equally.
@@
  • Linear Pricing. Under IIC, the cloud server determines a unit price $C$ for the client’s effort $e$ [bolton2004contract].
  • Conventional Federated Learning (CFL). FL methods integrating conventional incentive mechanisms (e.g., [wang2022infedge, sun2021pain, ding2023joint]), which do not consider CLPs, are used to compare with R3T proof of concept.
@@ IV-B Contract Feasibility
Impacts of different unit effort costs. In Figs. 4(4(a))-(4(f)), under CIC and IIC in CLPs [^9] We focus on one global training round in CLPs here, but analogous trends observed in non-CLPs., the client’s fixed salary, bonus, effort, total reward, utility, and cloud utility have shown decreasing trends per contract item along with the increase of unit effort cost $\delta$, where we set $\delta=\{0.2,0.6,1\}$. In other words, clients are not willing to join and contribute efforts when $\delta$ becomes large. For example, given the tenth contract item under the IIC cases, the values of the client’s effort units are around {9.60, 3.20, 1.92} corresponding to $\delta=\{0.2,0.6,1\}$. Besides, under CIC, where the cloud server possesses complete client information, the cloud can fine-tune and optimize its contract items to incentivize clients to join early and contribute more effort. This results in higher cloud utility while offering clients higher fixed salaries and bonuses, compared to IIC (e.g., 232.28 vs. 211.13, 5.96 vs. 1.01, and 119.21 vs. 18.47 given the tenth contract item with $\delta=0.2$, respectively). In addition, due to being aware of client information, the cloud server exploits the client’s efforts as much as possible while guaranteeing IR constraints (cf. Theorem 1) in CIC, leading client utilities over contract items equal to zero (cf., Fig. 4(4(e))).
@@
Impacts of different cloud budgets. Figs. 5(5(a))-(5(b)) show the effect of varying cloud budgets $P$ on the cloud utility contributed by type-$k$ clients (i.e., $k$-th contract items) in a global training round under both CIC and IIC in CLPs. As $P$ increases, more clients can be incentivized to participate in and allocate efforts, resulting in greater total cloud utility over all joined clients. For instance, at $P=60$, the total cloud utility of all type-$k$ clients reaches approximately 375 in CIC and 312 in IIC, compared to 266 in CIC and 311 in IIC at $P=30$, respectively. Especially at low budgets (i.e., $P=\{10,20,30\}$), the total cloud utility under IIC exceeds that under CIC. This is supported by Assumption 1 and the fact that the cloud server is aware of clients’ types in CIC and only the distribution of clients’ types in IIC, thereby different optimal contract designs. Consequently, in CIC, the cloud server may adopt a greedy allocation strategy to exhaust efforts from higher-type clients (cf. Theorem 1), but due to a limited budget, it leads to fewer participating clients and thus lower total cloud utility compared with that of IIC. Conversely, when the budget is huge enough (i.e., $P\geq 40$), the total cloud utility in CIC is far larger than that in IIC (i.e., 375 versus 311 at $P=60$).
@@ IV-D Proof of Concept of R3T
Convergence rate. As illustrated in Fig. 12, R3T demonstrates its efficiency by surpassing CFL in improving convergence rate under non-IID settings for both CIFAR-10 and Fashion-MNIST datasets. Particularly, R3T achieves a significant increase in performance during the initial FL training phase. This improvement is attributed to its strategy of attracting the highest-quality clients by offering an optimal set of contract items, which includes timely and substantial rewards, early in the training process. For example, R3T tends to reach the final target accuracy in just 20 rounds, compared with approximately 61 rounds for CFL on the datasets.
Long-term performance. For different values of the parameter of Dirichlet distribution $\alpha=\{0.4,0.7\}$ in both datasets, both R3T and CFL maintain stable test accuracies after convergence, even though R3T selects a smaller number of participating clients (cf., in line 1 of Algorithm 1). Under a highly non-IID setting (i.e., $\alpha=0.1$), R3T demonstrates more sustainable performance compared to CFL. In CFL, the practice of treating clients’ contributions and efforts equally across all training rounds, especially under a significant quantity and label distribution skew, detrimentally affects the long-term performance of the FL model, regardless of subsequent efforts. Besides, although R3T shows a slight drop in accuracy after CLPs due to the reduced number of selected clients, the overall test accuracy remains stable around the highest bound. These observations highlight the significant impact of attracting high-quality clients early, saving around 200%-300% in the total training time and 5.2%-47.6% in the total number of clients to achieve a close targeted learning performance.
V Conclusion
In this work, we studied R3T, a time-aware incentive framework for FL implemented on smart contracts. By characterizing the interactions between the client time and system capabilities, efforts, joining time, and their impact on cloud utility, we proposed a contract design that incorporates the temporal factor to derive optimal incentive strategies for the cloud. R3T transparently allocates the right reward at the right time, effectively attracting the highest-quality contributions during the early training phase, which is a critical determinant of final model performance under information symmetry and asymmetry. Simulation and proof-of-concept experimental results demonstrate the superiority of R3T over incentive benchmark mechanisms in cloud utility and economic efficiency. Furthermore, R3T reduces convergence time and the requisite number of participating clients, while achieving comparable model performance.